OpenStudy (anonymous):

sinθ=-9/10, θ is in quad IV, find sin and cos of 2θ

OpenStudy (loser66):

why do you post it twice?

OpenStudy (anonymous):

i didn't mean to

OpenStudy (loser66):

\(cos 2\theta = 1-2sin^2 \theta\) just plug the number in to get cos \(2\theta\)

OpenStudy (loser66):

about sin \(2\theta\) there are 2 easiest way to figure out, 1/ algebraic way 2/geometric way pick one

OpenStudy (loser66):

However, I guide you only, I don't work for you and I don't force you study. If you don't want to study, it's FINE to me

OpenStudy (anonymous):

I am working on it. . .

OpenStudy (anonymous):